# Sparse but Wrong: Incorrect L0 Leads to Incorrect Features in Sparse Autoencoders

### v1

@@ 4 LLM experiments
Next, we train a series of BatchTopK SAEs (Bussmann et al., 2024) on Gemma-2-2b (Team et al., 2024) layer 12 with different L0 values and calculate $s_{n}^{\text{dec}}$ for these SAEs. Each SAE is trained on 500M tokens from the Pile (Gao et al., 2020) using SAELens (Bloom et al., 2024) with a learning rate of $3e^{-4}$. During training, we ensure that the decoder remains normalized with $W_{dec}||_{2}=1$ so $s_{n}^{\text{dec}}$ calculations use the same scale for every latent. We train an SAE at each L0 from $10,20,30,\dots,250$. After L0=250, due to compute costs, we increase the spacing of the L0 of the SAEs up to L0=2500.
Figure 3: Nth decoder projection vs SAE L0 for N=700 (left) and N=10k (right) on Gemma-2-2b layer 12 SAEs with 32k width. Here as well, we see that $s_{n}^{\text{dec}}$ is minimized at the same point regardless of the choice of N, but the slope and shape of the $s_{n}^{\text{dec}}$ curve varies depending on the choice of N. These plots imply that the correct L0 for these SAEs is around 200-250, much higher than typically used.
We plot $s_{n}^{\text{dec}}$ vs L0 for Gemma-2-2b SAEs in Figure 3 for N=700 and N=10k (more plots in Appendix A.7). We see the same pattern for these SAEs as we saw in our toy model SAEs, namely that the metric is minimized at roughly the same L0, 200-250, despite the choice of N. As in the toy model SAEs case, using a higher N value results in a shallower slope leading up the the minimum $s_{n}^{\text{dec}}$ L0.
We test these SAEs using the sparse probing benchmark from Kantamneni et al. (2025), consisting of over 100 sparse probing tasks. We use the normal setting from this benchmark with k=1 and k=16 sparse probes. Results are shown in Figure 4. The results closely match what the $s_{n}^{\text{dec}}$ metric predicts, showing that L0 around 200-250 yields the highest results on the sparse probing benchmark.
Figure 4: Sparse probing F1 score vs L0 for Gemma-2-2b layer 12 SAEs. Results are shown for k=1 (left) and k=16 (right). Both results are maximized at L0 around 200-250, matching our $s_{n}^{\text{dec}}$ findings.
@@ 5 Related work
Chanin et al. (2025) explores feature hedging, showing SAEs mix correlated features into latents if the SAE is too narrow. We consider our work a version of feature hedging due to low L0. Till (2024) shows SAEs may increase sparsity by inventing features. Chanin et al. (2024) discuss the problem of feature absorption, where SAEs can improve their sparsity score by mixing hierarchical features together. Engels et al. (2024) investigates SAE errors and finds that SAE error may be pathological and non-linear.
@@ 6 Discussion
While most practitioners of SAEs understand that having too high of a L0 is problematic (at some point the sparse autoencoder is not sparse), our work shows that having too low of L0 is perhaps even worse. Our work has several important implications for the field. First, the L0 used by most SAEs is lower than it ideally should be, as a cursory search of open source SAEs on Neuronpedia (Lin, 2023) shows L0 less than 100 and even less than 50 is very common even for SAEs trained on large models (see Appendix A.5). Our work further shows that the framing of a sparsity–reconstruction tradeoff, as commonly discussed by most SAE papers (Cunningham et al., 2024; Gao et al., 2024; Rajamanoharan et al., 2024), is a misleading metric: when L0 is too low, an SAE with a correct dictionary achieves worse reconstruction than an incorrect SAE that mixes components of correlated features (see Appendix A.2).
While we expect that the correct L0 for a given SAE depends on the dataset used to train the SAE, we are also excited that our analysis may give us hints as to the true number of firing features in superposition in any given LLM activation.
The next step in our work is use our $s_{n}^{\text{dec}}$ metric to located the correct L0 for an SAE during training. We discuss one such technique in Appendix A.4. We also hope to estimate how the correct L0 varies with SAE and model size and layer, and investigate more SAE architectures.
Acknowledgments and Disclosure of Funding
David Chanin was supported thanks to EPSRC EP/S021566/1 and the Machine Learning Alignment and Theory Scholars (MATS) program.
Appendix A Technical Appendices and Supplementary Material
A.1 Toy model details
@@
A.2 Sparsity–reconstruction tradeoff in toy models
It is common in previous SAE work to discuss a sparsity–reconstruction tradeoff[Cunningham et al., 2024, Gao et al., 2024, Rajamanoharan et al., 2024], where the assumption is that having better reconstruction at a given sparsity is inherently better, and is a sign the SAE is correct. As we discussed in Section 3.1, when the L0 of the SAE is lower than optimal, the SAE can find ways to “cheat” and get a better MSE score by mixing components of correlated features together. This results in an SAE where the latents are no longer monosemantic, and do not track ground-truth features.
Figure 6: Sparsity ($x$) vs reconstruction ($y$) for learned SAEs and a ground-truth correct SAE at a given L0 in our toy model. When L0 is lower than the true L0 of the toy model, the learned SAE gets a higher reconstruction score than the ground-truth correct SAE, highlighting the problems with optimizing for reconstruction as a proxy for the quality of the SAE.
We explore the sparsity–reconstruction tradeoff in SAEs trained on our toy model at various L0s. Since we have the ground-truth correct features in our toy model, we also construct a ground-truth SAE that perfectly represents these true features. We manually set the L0 of the ground truth SAE while leaving the encoder and decoder fixed at the correct features. We plot the variance explained vs L0 in Figure 6 for the learned and ground-truth SAEs. We see that when the L0 of SAE is lower than the true L0 of the toy model, the ground-truth correct SAE actually scores worse on reconstruction than the incorrect learned SAE! This highlights the problems with optimizing reconstruction as a proxy for SAE quality.
(a) Learned SAE L0 = 1
(b) Learned SAE L0 = 2
(c) Ground truth SAE
Figure 7: SAE decoder cosine similarity with true features for the learned SAEs with L0=1 and L0=2, compared with the ground-truth correct SAE. The learned SAEs score more than twice as a high as the ground truth SAE on variance explained when L0=1 and L0=2, despite having extremely corrupted, polysemantic latents.
We show the cosine similarity of the SAE decoder latents with the ground truth features for the SAEs learned with L0=1 and L0=2 compared with the ground-truth SAE in Figure 7(c). Both these SAEs outperform the ground-truth SAE on variance explained by over 2x despite learning horribly polysemantic latents bearing little resemblance to the underlying true features of the model.
A.3 Transitioning L0 during training
@@
A.4 Automatically finding the correct L0 during training
@@
A natural next step of our finding that the correct L0 occurs when Nth decoder projection, $s_{n}^{\text{dec}}$, metric is minimized is to use this to find the correct L0 automatically during training. This is a meta-learning task, as the L0 is a hyperparameter of the training process. We find there are several challenges to directly using $s_{n}^{\text{dec}}$ as an optimization target:
@@
  • Small gradients directly above correct L0 In our plots of $s_{n}^{\text{dec}}$ from both toy models and Gemma-2-2b, we find that the metric is relatively flat in a region start at the correct L0 and extending to higher L0 values. We thus need a way to traverse this flat region and stop once the metric starts to increase again.
@@
  • The impact of changing L0 is delayed We find that it takes many steps after changing L0 for $s_{n}^{\text{dec}}$ to also change, meaning it is easy to overshoot the target L0 or oscillate back and forth.
@@
  • Noise during training We find that while $s_{n}^{\text{dec}}$ shows clear trends after training for many steps, it can be noisy on each training sample. So our optimization needs to be robust to this noise.
Taking these requirements into account, we present an optimization procedure to find the L0 that minimizes $s_{n}^{\text{dec}}$ automatically during training. We first estimate the gradient of $s_{n}^{\text{dec}}$, hereafter referred to as to as the metric, $m$, with respect to L0, $\nicefrac{{dm}}{{dL0}}$. We first define an evaluation step $t$ as a set number of training steps (we evaluate every 100 training steps). At $t$ we change L0 by $\delta_{L_{0}}$. At the next evaluation step, $t+1$, we evaluate $m$. We use a sliding average of $s_{n}^{\text{dec}}$ over the past 10 training steps to calculate $m$ to help account for noise. We the estimate $\nicefrac{{dm}}{{dL0}}$ as:
@@
We then provide this gradient to the Adam optimizer [Kingma and Ba, 2014] with default settings, and allow it to change the L0 parameter.
@@
We find that this optimization strategy works very well in toy models, but requires a lot of hyperparameter tuning to work in real LLMs. The starting L0, $n$ for $s_{n}^{\text{dec}}$, $b$, learning rate for the Adam optimizer, and min and max $\delta_{L_{0}}$ values all have a big impact on how fast and how aggressively the optimization works. The slop of $m$ around the correct L0 is shallow, so it is easy to overshoot. We also find that different values of $n$ take more or less time to converge during training. We expect it is possible to further simplify and improve this process in future work.
A.5 L0 of open-source SAEs on Neuronpedia
We analyze common open-source SAEs as provided by Neuronpedia [Lin, 2023] and SAELens [Bloom et al., 2024]. We include all SAEs cross-listed in both SAELens and Neuronpedia with an L0 reported in SAELens. We show the results as a histogram in Figure 9. Our analysis shows that for layer 12 of Gemma-2-2b, the correct L0 should be around 200-250. However, we find that most open-source SAEs have L0 below 100, much lower than our analysis expects to be ideal.

### v2

@@ 4 LLM experiments
We train a series of BatchTopK SAEs (Bussmann et al., 2024) with $h=32768$ on Gemma-2-2b (Team et al., 2024) and Llama-3.2-1b (Dubey et al., 2024) varying L0 and calculate $s_{n}^{\text{\text{dec}}}$. Each SAE is trained on 500M tokens from the Pile (Gao et al., 2020) using SAELens (Bloom et al., 2024). We also calculate k-sparse probing performance for these SAEs using the benchmark from Kantamneni et al. (2025), consisting of over 100 sparse probing tasks. Results are shown in Figure 9 for $n=16000$. We find that using $n$ near $h/2$ seems to achieve best results in our LLM experiments.
Figure 9: $n$ = 16k decoder projection vs SAE L0 and K-sparse probing F1 vs L0 with 3 seeds per L0. (left) Gemma-2-2b layer 5 BatchTopK results. (right) Llama-3.2-1b layer 7 BatchTopK SAEs. In both cases, peak sparse probing performance occurs in the elbow just before $s_{n}^{\text{\text{dec}}}$ jumps due to low L0, although the shapes of the $s_{n}^{\text{\text{dec}}}$ plots vary at high L0.
The Llama SAE $s_{n}^{\text{dec}}$ plot looks very similar to the toy model, with a clear minimum point. The Gemma-2-2b layer 5 SAEs also show a sharp increase in $s_{n}^{\text{dec}}$ at low L0 as we saw in toy models, but has a long shallow region with the global minimum actually appearing in that shallow region. In both cases, the “elbow” in the $s_{n}^{\text{dec}}$ plots just before the jump due to low L0 is around L0 200, and this also corresponds to peak sparse probing performance. More plots and analysis of $s_{n}^{\text{dec}}$ curves are shown in Appendix A.9.
Figure 10: Gemma-2-2b layer 12, with (left) $n$th decoder projection and K-sparse probing F1 for BatchTopK and JumpReLU SAEs, and (right) normalized decoder projection histograms for BatchTopK SAEs. The histograms are truncated to -20 and 40 to highlight projections near the origin.
4.1 JumpReLU vs BatchTopK SAEs
We next explore how JumpReLU and BatchTopK SAEs compare with $n$th decoder projection plots. We train a suite of SAEs on 1B tokens on Gemma-2-2b layer 12. We plot $s_{n}^{\text{dec}}$ for a range N values as well as k-sparse probing results for JumpReLU and BatchTopK SAEs in Figure 10 (left).
JumpReLU and BatchTopK SAEs behave similarly at low L0, with the high $s_{n}^{\text{dec}}$ at low L0 corresponding to poor sparse-probing performance. However, we see notable differences at high L0. The BatchTopK SAEs have a global $s_{n}^{\text{dec}}$ minimum around 200-250 no matter the N value, but JumpReLU SAEs only have a clear $s_{n}^{\text{dec}}$ minimum at 200-250 when $n=16000$ (about $h/2$). As we saw in Figure 9 as well, using the “elbow” of the plots just before $s_{n}^{\text{dec}}$ jumps due to low L0 seems to correspond to peak k-sparse probing performance.
For JumpReLU SAEs, we see that $s_{n}^{\text{dec}}$ does not rise much at high L0, and indeed, JumpReLU SAEs also perform much better than BatchTopK SAEs at sparse probing when L0 is high. We suspect this is due to JumpReLU SAEs being able to “stick” near the correct threshold per latent like we saw in our toy models section. We investigate the differences in learned SAEs between JumpReLU and BatchTopK further in Appendix A.10.
4.2 Can L0 be both too low and too high simultaneously?
In Figure 10 (right), we plot decoder histogram projection plots for BatchTopK SAEs on Gemma-2-2b layer 12 with L0 10, 200, 750, and 2000. As we expected from Figure 6, when L0 is very low (10) or very high (2000), we see a wide gaussian around 0, indicating that decoder latents are mixing correlated features together. At L0=200, we see a much more narrow distribution around 0, as we expect when near the correct L0. However, at L0=750, we see an interesting phenomenon, where there is an even narrower distribution than at L0=200, but also a large hump starting at projection above 10 (more visible in the log plot).
We suspect this indicates at L0=750, some latents become more monosemantic while other latents mix underlying features becoming less monosemantic. This likely means that the L0 is too high for some latents while simultaneously being too low for other latents. There is no reason why every latent has the same optimal L0 threshold, so there is likely a range of L0s where some latents are firing more than they ideally should while other latents are firing less than they ideally should. We also suspect this is part of why JumpReLU SAEs seem to perform much better at high L0, since JumpReLU SAEs can adjust firing threshold per-latent while BatchTopK SAEs cannot.
@@ 5 Related work
Chanin et al. (2025) explores feature hedging, showing SAEs mix correlated features into latents if the SAE is too narrow. We consider our work a version of feature hedging due to low L0. Till (2024) shows SAEs may increase sparsity by inventing features. Chanin et al. (2024) discuss the problem of feature absorption, where SAEs can improve their sparsity score by mixing hierarchical features together. Engels et al. (2024) investigates SAE errors and finds that SAE error may be pathological and non-linear. Engels et al. (2025) find that not all underlying LLM features themselves are linear, demonstrating circular embeddings of some concepts. Wu et al. (2025) and Kantamneni et al. (2025) both investigate empirical SAE performance, finding SAEs underperform relative to supervised baselines, but do not offer theoretical explanations as to why SAEs underperform.
@@ 6 Discussion
While most practitioners of SAEs understand that having too high L0 is problematic, our work shows that having too low of L0 is perhaps even worse. Our work has several important implications for the field. First, the L0 used by most SAEs is lower than it ideally should be, as a cursory search of open source SAEs on Neuronpedia (Lin, 2023) shows L0 less than 100 is very common even for SAEs trained on large models (see Appendix A.7). We further show that the sparsity–reconstruction tradeoff, as commonly discussed by most SAE papers (Cunningham et al., 2024; Gao et al., 2024; Rajamanoharan et al., 2024), is misleading: when L0 is too low, an SAE with a correct dictionary achieves worse reconstruction than an incorrect SAE that mixes correlated features.
We presented a metric based on the correlation between the SAE decoder and input activations, $s_{n}^{\text{\text{dec}}}$, that can give us hints about the correct L0 for a given SAE. However, we do not view this as a perfect guide. As we saw in our results, while low L0 SAEs consistently have very high $s_{n}^{\text{\text{dec}}}$, the metric is not always high at high L0 (although it usually is). Still, we feel that this metric is a useful guide to avoid L0 that is clearly too low, and we hope this investigation into correlation-based SAE quality metrics can be built on further in future work.
While our metric currently requires training a sweep over L0 to optimize, we are hopeful that it may be possible to optimize this metric automatically during training (steps towards this are discussed in Appendix A.6). We further encourage anyone implementing ideas from this work to also experiment with our alternative metric, decoder pairwise cosine similarity, discussed in Appendix A.5.
Acknowledgments
David Chanin was supported thanks to EPSRC EP/S021566/1 and the Machine Learning Alignment and Theory Scholars (MATS) program. We are grateful to Henning Bartsch for feedback during the project.
Appendix A Appendix
A.1 SAE training architecture definitions
In this work we focus on JumpReLU (Conerly et al., 2025; Rajamanoharan et al., 2024) and BatchTopK (Bussmann et al., 2024) SAEs. For BatchTopK SAEs, there is no sparsity penalty as sparsity is enforced by the BatchTopK function. The auxiliary loss $\mathcal{L}_{P}$ for BatchTopK is as follows, where $e$ is the SAE training error residual, and $\hat{e}$ is a reconstruction using the top $k_{\text{aux}}$ dead latents (meaning the latents have not fired in more than $n_{\text{dead}}$ steps during training).
$\mathcal{L}_{p}=\lVert e-\hat{e}\rVert_{2}^{2}$
We follow the JumpReLU training setup from Conerly et al. (2025), which involves both a sparsity loss $\mathcal{L}_{s}$ and a pre-activation loss for reviving dead latents, $\mathcal{L}_{p}$. $\mathcal{L}_{s}$ is defined as below, where $c$ is a scalar scaling factor:
$\mathcal{L}_{s}=\sum_{i}\text{tanh}(c*|a_{i}|\lVert\mathbf{W}_{\text{dec},i}\rVert_{2})$
The pre-activation loss $\mathcal{L}_{p}$ adds a small penalty for all dead features, where $a_{\text{pre}}$ refers to the pre-activation of the SAE passed into the JumpReLU:
$\mathcal{L}_{p}=\sum_{i}\text{ReLU}(\tau_{i}-a_{\text{pre},i})\lVert\mathbf{W}_{\text{dec},i}\rVert_{2}$
The JumpReLU defines a pseudo-gradient relative to the threshold $\tau$ as follows, where $\epsilon$ is the bandwidth of the estimator:
$\frac{\partial\text{JumpReLU}(x,\tau)}{\partial\tau}=\begin{cases}-\frac{\tau}{\epsilon}&\text{if }-\frac{1}{2}<\frac{x-\tau}{\epsilon}<\frac{1}{2}\\ 0&\text{otherwise}\end{cases}$
A.2 Toy model SAE training details
We train on 15M samples with a batch size of 1024 for all toy model experiments, and a learning rate of 3e-4. We do not use any learning rate warm-up or decay. For all SAE latents vs true feature cosine similarity plots, we re-arrange the SAE latents so the latent indices match the feature indices in the plots, as this makes interpreting the plots easier without any loss of generality.
For the large toy model experiments in Section 3.2, we use a randomly generated correlation matrix and linearly decreasing feature firing probabilities, both shown in Figure 11.
@@ A.2 Toy model SAE training details
A.3 LLM SAE training details
For BatchTopK SAEs, we ensure that the decoder remains normalized with $\mathbf{W}_{\text{dec}}||_{2}=1$ so $s_{n}^{\text{\text{dec}}}$ calculations use the same scale for every latent. We use a learning rate of $3e^{-4}$ with no warmup or decay.
For JumpReLU SAEs, we broadly follow the training procedure laid out by Conerly et al. (2025). However, we do not apply learning rate decay, and only warm $\lambda_{s}$ for 100M tokens to avoid the sparsity penalty changing throughout the majority of training. We use a learning rate of $2e^{-4}$, $c=4$, $\lambda_{p}=3e-6$ and bandwidth $\epsilon=2.0$ as recommended by (Conerly et al., 2025).
A.4 Transitioning L0 during training
@@
A.5 Alternative metric: decoder pairwise cosine similarity
We find that $n$th decoder projection is not the only metric that works to gauge the extent to which features are being mixed across SAE latents. Simply taking the mean cosine similarity of each decoder latent with every other decoder latent seems to work as well. We call this metric decoder pairwise cosine similarity, $c$, and define it as below:
$c=\frac{1}{\binom{h}{2}}\sum_{i=1}^{h-1}\sum_{j=i+1}^{h}|\cos(\mathbf{W}_{\text{dec},i},\mathbf{W}_{\text{dec},j})|$ (6)
where $\binom{h}{2}=\frac{h(h-1)}{2}$ is the total number of distinct pairs.
The intuition behind this metric is that if SAE decoder latents are mixing lots of positive and negative components of correlated and anti-correlated features, then each SAE latent should become less orthogonal to each other SAE latent, as many latents will likely mix together similar features. This should mean that the absolute value of the cosine similarity between arbitrary latents should also increase the worse this mixing becomes.
A.5.1 Toy model results
We calculate pairwise calculate similarity $c$ for each of the BatchTopK SAEs we trained on toy models from Section 3.5. Results are shown in Figure 13. We see that pairwise cosine similarity is minimized at the true L0.
Figure 13: Decoder pairwise cosine similarity evaluated on 5 seeds of toy model SAEs. The true L0 is indicated with a dotted line at 11. The shaded area is 1 stdev. The metric is minimized at the true L0.
A.5.2 LLM SAE results
Next, we calculate pairwise cosine similarity for each LLM SAE we evaluated in the paper along with k=16 sparse probing results. Gemma-2-2b layer 5 and Llama-3.2-1b layer 7 results are shown in Figure 14. The results roughly match what we saw with $s_{n}^{\text{dec}}$, being minimized at perhaps a slightly lower but still similar L0.
Figure 14: Decoder pairwise cosine similarity metric and k=16 sparse probing results for BatchTopK SAEs trained on Gemma-2-2b layer 5 (left) and Llama-3.2-1b layer 7 (right). The metric is roughly minimized near peak sparse-probing performance. The shaded area is 1 stdev.
Figure 15: Decoder pairwise cosine similarity metric and sparse probing results for BatchTopK and JumpReLU SAEs trained on Gemma-2-2b layer 12. The metric seems to align less well with k=16 sparse probing results (left), but aligns reasonably well with k=1 sparse probing results (right). The pairwise cosine similarity plot is repeated on both sides to make a vertical comparison with the k-sparse probing results easier.
BatchTopK and JumpReLU results for Gemma-2-2b layer 12 are shown in Figure 15. Interestingly, the BatchTopK results seem to be minimized at a slightly earlier L0 than the JumpReLU results under this metric, which is different than the $n$th decoder projection results from Section 4, where we saw both minimized at roughly the same point. We note that the decoder pairwise cosine similarity metric seems to align better with k=1 sparse probing results, while $n$th decoder projection seems to align better with k=16 sparse probing results. We do not know if this is a meaningful difference between these metrics. It is possible that, as we saw in Section 4.2, there is a range of L0 values that can be too high for some latents while too low for other latents in BatchTopK SAEs, and perhaps the difference between metrics (and k=1 vs k=16 sparse probing results) relates to differences in sensitivity to this mixture of too high / too low latents. Regardless, in all cases, when L0 is low both metrics show large values, indicating high polysemanticity in SAE latents.
A.5.3 Which metric is better?
We choose to focus on $n$th decoder projection in this paper since we feel it allows for more insightful interpretations of the decoder projection plots (e.g. Section 4.2). However, we do not feel that either metric is inherently better or worse than the other, and we do not have a strong reason to pick either metric. We expect that when an SAE is near the correct L0, there are likely many indicators that all should point to similar results. We thus recommend that practitioners try both metrics to see which results in easier interpretation for their use-case.
A.6 Automatically finding the correct L0 during training
@@
A natural next step of our finding that the correct L0 occurs when Nth decoder projection, $s_{n}^{\text{\text{dec}}}$, metric is minimized is to use this to find the correct L0 automatically during training. This is a meta-learning task, as the L0 is a hyperparameter of the training process. We find there are several challenges to directly using $s_{n}^{\text{\text{dec}}}$ as an optimization target:
@@
  • Small gradients directly above correct L0 In our plots of $s_{n}^{\text{\text{dec}}}$ from both toy models and Gemma-2-2b, we find that the metric is relatively flat in a region start at the correct L0 and extending to higher L0 values. We thus need a way to traverse this flat region and stop once the metric starts to increase again.
@@
  • The impact of changing L0 is delayed We find that it takes many steps after changing L0 for $s_{n}^{\text{\text{dec}}}$ to also change, meaning it is easy to overshoot the target L0 or oscillate back and forth.
@@
  • Noise during training We find that while $s_{n}^{\text{\text{dec}}}$ shows clear trends after training for many steps, it can be noisy on each training sample. So our optimization needs to be robust to this noise.
Taking these requirements into account, we present an optimization procedure to find the L0 that minimizes $s_{n}^{\text{\text{dec}}}$ automatically during training. We first estimate the gradient of $s_{n}^{\text{\text{dec}}}$, hereafter referred to as to as the metric, $m$, with respect to L0, $\nicefrac{{dm}}{{dL0}}$. We first define an evaluation step $t$ as a set number of training steps (we evaluate every 100 training steps). At $t$ we change L0 by $\delta_{L_{0}}$. At the next evaluation step, $t+1$, we evaluate $m$. We use a sliding average of $s_{n}^{\text{\text{dec}}}$ over the past 10 training steps to calculate $m$ to help account for noise. We the estimate $\nicefrac{{dm}}{{dL0}}$ as:
@@
We then provide this gradient to the Adam optimizer (Kingma & Ba, 2014) with default settings, and allow it to change the L0 parameter.
@@
We find that this optimization strategy works very well in toy models, but requires a lot of hyperparameter tuning to work in real LLMs, limiting its utility. The starting L0, $n$ for $s_{n}^{\text{\text{dec}}}$, $b$, learning rate for the Adam optimizer, and min and max $\delta_{L_{0}}$ values all have a big impact on how fast and how aggressively the optimization works. The slope of $m$ around the correct L0 is shallow, so it is easy to overshoot. We also find that different values of $n$ take more or less time to converge during training. We expect it is possible to further simplify and improve this process in future work.
A.7 L0 of open-source SAEs on Neuronpedia
We analyze common open-source SAEs as provided by Neuronpedia (Lin, 2023) and SAELens (Bloom et al., 2024). We include all SAEs cross-listed in both SAELens and Neuronpedia with an L0 reported in SAELens. We show the results as a histogram in Figure 16. Our analysis shows that for layer 12 of Gemma-2-2b, the correct L0 should be around 200-250. However, we find that most open-source SAEs have L0 below 100, much lower than our analysis expects to be ideal.